If m is the mass of the earth and r is its radius,the ratio of the gravitational Acceleration and the gravitational constant at its surface is what?

Grade:9

1 Answers

Utsav Basu
6 years ago
Let mo be the mass of any object and g represent the gravitational acceleration on the surface of the earth.
 
Thus, we can write,
     mog=G(mmo)/r2
or, g=Gm/r2
or, g/G=m/r2
 
So, the ratio of the gravitational Acceleration and the gravitational constant at its surface is m/r2.